OpenStudy (anonymous):

can someone show me how to find the zeros... f(x) =x^3+4x^2+4x. FIND the zeros of f.

OpenStudy (anonymous):

set it = 0 so you can factoooorrrr

OpenStudy (anonymous):

and you can factor out an x from the get go!

OpenStudy (anonymous):

0= x(x²+4x-4)

OpenStudy (anonymous):

right x(x^2+4x+4)=0 x(x+2)^2=0

OpenStudy (anonymous):

0=x(x +2) (x-2) x = 0 x = 2 x= -2 these are the zeros.

OpenStudy (anonymous):

x=0, and x+2=0 is x=-2

OpenStudy (saifoo.khan):

f is written just beside x.

OpenStudy (anonymous):

lol yes, correct saifooooo.

OpenStudy (anonymous):

^im confused... >.<

OpenStudy (saifoo.khan):

are u?

OpenStudy (anonymous):

NEVER MIND. got it! haha

OpenStudy (anonymous):

yayyyyyyyyyyy

OpenStudy (anonymous):

yo

OpenStudy (anonymous):

sup

OpenStudy (anonymous):

im not so good at graphs..so.. f(x)=1-3^x the graph of f is reflected over the line y=x. what is the name of function of the new graph?